OpenStudy (anonymous):

At Ron's Roller Rink, the number of customers has been decreasing at steady rate of 5% per year. if there were 900 skaters per week in 2000, what is a good estimate for the number skaters per week in 2006 ?

OpenStudy (rsadhvika):

900(1-.05)^(2006-2000) = 900(1-.05)^(6) = 900(.95)^6

OpenStudy (rsadhvika):

the decay forms a geometric series